OpenStudy (anonymous):

use partial derivatives to find dy/dx if y=f(x) is determined implicitly by the given equation. 2x^3+x^2y+y^3=1

OpenStudy (anonymous):

\[2x^3+x^2y+3=1\]

OpenStudy (amistre64):

you dont use partial for implicits

OpenStudy (anonymous):

thats what the directions on the hw say lol

OpenStudy (amistre64):

implicits imply that y changes as a result of x partials are when y changes independant of x

OpenStudy (anonymous):

i don't know what you want me to say lol, this is directly from the book..

OpenStudy (amistre64):

then get a new book :) there is no way I see that this can be answered as is with any degree of intelligence ...

OpenStudy (amistre64):

as is, it would simply be an implicit differentation

OpenStudy (anonymous):

the answer is : \[- (6x^2+2xy) / (x^2+3y^2)\] if that sheds any light? cuz i'm completely lost.. this chapter is on the chain rule?

OpenStudy (amistre64):

6x^2+2xy+x^2y'+3y^2 y' = 0

OpenStudy (amistre64):

yep, its just an implicit ....

OpenStudy (amistre64):

move the non y's to the right; then factor out your y' and divide off the slag

OpenStudy (anonymous):

you've lost me ^.^

OpenStudy (amistre64):

do you know how to do implicit differentation?

OpenStudy (anonymous):

no :(

OpenStudy (amistre64):

do you know how to do regular differentation?

OpenStudy (anonymous):

yes!

OpenStudy (amistre64):

then you know implicit, youve just not been taught that you know it yet

OpenStudy (amistre64):

your used to getting rid of the x' bit

OpenStudy (amistre64):

for example; tell me, what is the derivative of 2x?

OpenStudy (anonymous):

2

OpenStudy (amistre64):

nope, it 2x' i never said "with respect to x"

OpenStudy (amistre64):

what is the derivative of 5y^2 ?

OpenStudy (anonymous):

5y^2' ? lol

OpenStudy (amistre64):

10y y' ...

OpenStudy (amistre64):

whats the derivative of sin(2t) ?

OpenStudy (anonymous):

now you've lost me lol 2sin(2t)'?!

OpenStudy (amistre64):

you learn from the start how to derive things with respect to x ... \[\frac{d}{dx}4x^3\to \frac{dx}{dx}12x^2\] since dx/dx = 1 you think its not there

OpenStudy (amistre64):

2cos(2t) t'

OpenStudy (amistre64):

assume that we dont give a "with respect to .... \[\frac{d}{d*}4x^3\to\ \frac{dx}{d*}12x^2\] know we dont know what x' is equal to do we?

OpenStudy (amistre64):

\[\frac{d}{d*}4xy^2\to\ \frac{d4}{d*}xy^2+\frac{dx}{d*}4y^2+\frac{dy}{d*}4x2y\] \[x'\ 4y^2+y'\ 8xy\]

OpenStudy (amistre64):

if *=x , its simplifies to \[4y^2+y'\ 8xy\]

OpenStudy (amistre64):

all implicits are doing is not throwing out the derived bits

OpenStudy (amistre64):

product rule is still product rule power rule is still power rule and all the other rules youve learned are still all the other rules; nothing changes.

OpenStudy (anonymous):

so 2x^3 is 6x^2x' ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so if its asking for dy/dx, what is it asking me to derive? o.O

OpenStudy (amistre64):

yes; but derive it with the bits left in; then since this defines /dx all the x' s = 1

OpenStudy (amistre64):

all the y' s are just dy/dx

OpenStudy (amistre64):

\[x^2y\] derive it with power and product rule

OpenStudy (amistre64):

mostly product rule; but power when needed :)

OpenStudy (anonymous):

uhh...... x^2x'+2x'y'?!? o.O

OpenStudy (amistre64):

almost, keep your wits about you first split it into the product rule x^2' y + x^2 y' ; then derive the parts 2xx' y + x^2 y' ; and since this is /dx then x' = 1 2xy + x^2 y'

OpenStudy (amistre64):

derive this one now y^3

OpenStudy (anonymous):

3y^2y'?

OpenStudy (amistre64):

exactly

OpenStudy (amistre64):

and the constant after the equals sign derives to 0 like every good constant needs to

OpenStudy (amistre64):

put it all together and slove it out for y'

OpenStudy (anonymous):

2xy + x^2 y' why isn't there an x' in this?

OpenStudy (amistre64):

there was, but read the next line; since dx/dx = 1 we can simplify it

OpenStudy (anonymous):

oohhhh okay that makes sense! i think i get it now! thank you!!!!

OpenStudy (amistre64):

practice keeping all your derived bit in ; and then the variables that match the /d* part go to one :) good luck